# Failure of the Mott’s formula for the Thermopower in Carbon Nanotubes

### v1

@@ Abstract
The well-known Mott’s formula links the thermoelectric power characterised by Seebeck coefficient to conductivity. We calculate analytically the thermoelectric current and Seebeck coefficient in one-dimensional systems and show that, while the prediction of Mott’s formula is valid for Dirac fermions, it is misleading for the carriers having a parabolic dispersion. We apply the developed formalism to metallic single wall carbon nanotubes and obtain a non-trivial non-monotonic dependence of the Seebeck coefficient on the chemical potential. We emphasize that, in contrast to the Mott’s formula, the classical Kelvin’s formula that links thermoelectric power to the temperature derivative of the chemical potential is perfectly valid in carbon nanotubes in the ballistic regime. Interestingly, however, the Kelvin’s formula fails in two- and three-dimensional systems in the ballistic regime.
@@
$S_{K}=\frac{1}{e}\left(\frac{d\mu}{dT}\right).$ (2)
The alternative, kinetic approach to the theoretical description of the Seebeck effect was developed by Sir Nevil Mott in the second half of the XXth century [9]. The Mott’s formula relates the Seebeck coefficient to the conductivity of the system:
@@
where $\sigma$ is conductivity, $\mu$ is the chemical potential, $T$ is temperature. Here and further we assume $k_{B}=1.$ One can see that substituting the above expression for conductance to the Mott’s formula one obtains $S_{M}=0$ for any value of chemical potential except the close vicinity of the bottoms of the quantisation subband. Below we demonstrate that this simple conclusion fails and the Mott’s formula is unable to describe the thermoelectric phenomena in 1D systems in the ballistic regime. In contrast, the Kelvin’s formula remains valid in the ballistic regime in any one-dimensional (1D) system. Interestingly, however, the Kelvin’s formula turns out to be no more valid in two- and three-dimensional (2D and 3D) systems, in the ballistic regime.
The thermocurrent and thermoelectricity in the ballistic regime have been carefully studied theoretically in 1D, 2D and 3D systems [10, 11, 12]. The difference between these works and our present work is that we focus on the broken circuit geometry where no electric current is flowing through the system. We present an analytical theory of the Seebeck effect in metallic CNTs in the regime of a ballistic transport. We note that the Seebeck coefficient is also a direct measure of the entropy per particle, which makes it one of the most important characteristics of the statistics of quasiparticles in crystals [13]. We find that, in CNTs, it is dramatically dependent on the type of electronic dispersion. In the case of a parabolic dispersion, the Seebeck coefficient is a non-monotonic function of the chemical potential. Its magnitude decreases with the increase of the chemical potential, and its sign changes in the vicinity of the resonances of the chemical potentials and bottom energies of the electron and hole quantization subbands. In contrast,in the case of a linear dispersion that is observed in the vicinity of Dirac points in conducting CNTs the Seebeck coefficient is equal to zero. The sharp contrast between the behaviors of conductivity and Seebeck coefficient as functions of the chemical potential that may be efficiently controlled by an applied bias opens room for a variety of non-trivial effects governed by an interplay of currents induced by electric field and temperature gradient.
Recovering of the Kelvin’s formula for Seebeck coefficient in the ballistic regime for a 1D system. The definition of the Seebeck coefficient in the case of a broken electric circuit (${\cal J}=0$, where ${\cal J}$ is an electric current) reads
@@
Here and in what follows we assume the degeneration factor of quantum states $g=1$. The results can be easily generalized for any other value of $g$. The electronic velocity is defined by the general relation:
@@
No surprises, their product turns out to be constant $\nu(E)v(E)=\frac{1}{h}.$
@@
Coming back to the 1D case, one can rewrite the Kelvin’s formula in terms of the logarithmic derivative of DOS (see Supplemental Material, Sec. S1 [14]).
$S_{K}=\frac{1}{e}\left(\frac{\partial\mu}{\partial T}\right)=\frac{\pi^{2}}{3e}T\frac{\partial}{\partial\mu}\ln\nu(\mu).$ (11)
Conductance and the Mott’s formula for 1D two-subband system in ballistic regime. It is instructive to compare the exact expression for the Seebeck coefficient obtained above with an approximation that follows from the Mott’s formula. In order to derive it, we start from the generic expression for the electric current density in a one-dimensional channel in the ballistic regime (5).
@@
$\nu(E)=\frac{1}{2\pi}|\frac{\partial E}{\partial k}|^{-1}\left[1+\theta\left(E-E_{p}\right)\right],$ (12)
@@
This expression can be now substituted to the Mott’s formula (3). As a result we obtain:
@@
The specific case of a metallic CNT. Energy spectra of single-walled carbon nanotubes.
Figure 1: (a) A sketch of a zigzag CNT with chirality $(9,0)$. (b) The energy band structure of a quasi-metallic CNT according to the dispersion relation given by Eq. (15). Blue lines correspond to the energy levels with a quantum number $j=-1$ and $j=1$, green lines are for the energy levels with $j=0$. A small band gap due to the curvature cannot be seen at this energy scale.
Basic electronic properties of single-walled (dubbed originally as single-shell [15]) carbon nanotubes (CNTs) have been well-understood for over three decades [16] and depend entirely on their chirality $(m,n)$, where (as explained in detail in Ref. [17]) the numbers $(m,n)$ express the wrapping ‘chiral’ vector connecting two carbon sites along the nanotube circumference via the two unit vectors of the underlying hexagonal graphene lattice. A CNT with $n-m\neq 3p$, where $p$ is an integer, is a quasi-one-dimensional semiconductor with the band gap $E_{g}$ inversely proportional to the nanotube radius $R=a\sqrt{n^{2}+m^{2}+nm}/(2\pi)\approx 39.15\sqrt{n^{2}+m^{2}+nm}$ pm, where the graphene lattice constant $a=\sqrt{3}a_{C-C}$ with $a_{C-C}=1.42$ being a distance between neighbouring carbon atoms in graphene. Only armchair $(n,n)$ nanotubes are truly metallic (gapless) with the low-energy part of dispersion well-described by $E=\pm\hbar v_{F}|k-k_{0}|$, where $v_{F}\approx 9.8\times 10^{5}$ m/s is the Fermi velocity in graphene and $k_{0}=\pm\pi/(3a)$.
@@
For the two branches closest to the Dirac point ($j=0$), for the metallic and quasi-metallic CNTs $E_{0}=0$. The bottoms of their higher subbands are given by $E_{j}=j\hbar v_{F}/R$. Notably, the subbands with $|j|\geq 1$ are reasonably-well described by the effective mass approximation with $m_{j}=E_{j}/v_{F}^{2}$.
The specific case of a metallic CNT. A vanishing contribution to the Seebeck coefficient from linear dispersion subbands. For the linear dispersion subband DOS is constant and given by $\nu_{0}(\mu)=\frac{2}{\pi v_{F}}$, where $v_{F}$ is the sound velocity. As a result, the temperature derivative of the concentration of charge carriers turns zero. Hence the Dirac mode does not contribute to the Seebeck coefficient in agreement with both the Kelvin’s and Mott’s formulae. This statement remains valid also in the vicinity of the Dirac point due to the compensation of electron and hole contributions to the Seebeck effect (see Supplemental Material, Sec. S2 [14]).
Conflicting predictions of the Kelvin’s and Mott’s approaches for the contributions of parabolic subbands to the Seebeck coefficient.
Figure 2: DOS of a metallic CNT without the smearing effect (black line) and with smearing one for $\tau$=483.09 $eV^{-1}$ (blue line), $\tau$=38.65 $eV^{-1}$ (green line) and $\tau$=3.86 $eV^{-1}$ (red line).
@@
$\nu(\mu)=\sum\limits_{j=1}^{J_{\max}}\frac{\sqrt{2m_{j}}}{\pi}\frac{1}{\sqrt{\mu-E_{j}}},$ (16)
@@
Figure 3: The Seebeck coefficient for a metallic CNT as a function of chemical potential $\mu$ calculated for three temperatures T = 2.4 K (a), T = 77 K (b) and T = 300 K (c) according to Kelvin’s formula Eq. (17) (red lines) and Mott’s formula Eq. (14) (blue lines).
@@
The dependence of the Seebeck coefficient on the chemical potential calculated at different temperatures is shown in Fig. 3. One can notice the strong difference in the results obtained following the Kelvin’s approach (red lines) and Mott’s approach (blue lines). The Kelvin’s approach predicts the hyperbolic decay of the Seebeck coefficient away from the resonance between the chemical potential and the bottom of the subband. At the resonance point, the singularity and the change of sign of the Seebeck coefficient are observed. In contrast, the Mott’s formula predicts a finite value of the Seebeck coefficient at the bottom of subband, no sign change and a fast exponential decay away from the resonance. We note at this point that long ago the change of the sign of the Seebeck coefficient was identified as a signature of the topological phase transition which indeed takes place ones the chemical potential crosses the bottom of the next quantization subband [18, 19, 20, 21, 22, 23, 24]. The divergence of Seebeck coefficient in absence of scattering is caused by the singularity in DOS. It disappears once the smearing of the density of states is taken into account as we show in the next section.
Effect of smearing of DOS in parabolic subbands.        The Kelvin’s formula. In any realistic 1D system including CNTs, singularities of DOS are smeared due to a variety of factors: from finite size effects to fluctuations and non-linearities. Therefore, it is important to consider an effect of smearing of DOS on the Seebeck coefficient in the ballistic regime. The Green function of an electron in a single-wall CNT is
$G_{j}^{R}(E,p)=\frac{1}{E-p_{z}^{2}/2m_{j}-E_{j}+i/2\tau},$ (18)
where $\tau$ is the positive parameter responsible for the smearing effect and $p_{z}$ is the continuous momentum along the $z$-direction.
@@
$\nu^{(sm)}(E)=\sum\limits_{j=1}^{J_{\max}}\frac{\sqrt{m_{j}}}{\pi}\frac{\sqrt{\sqrt{\left(E-E_{j}\right)^{2}+\frac{1}{4\tau^{2}}}+\left(E-E_{j}\right)}}{\sqrt{\left(E-E_{j}\right)^{2}+\frac{1}{4\tau^{2}}}},$ (20)
@@
Now one can derive the exact expression for the contribution to the Seebeck coefficient of a CNT coming from parabolic energy subbands:
$S^{(sm)}_{K}(\mu)=\frac{\pi^{2}T}{3eJ_{\max}}\frac{\partial\ln\nu^{(sm)}(\mu)}{\partial\mu}=\frac{\pi^{2}T}{6eJ_{\max}}\frac{\sum\limits_{j=1}^{J_{\max}}\sqrt{m_{j}}\frac{\sqrt{\sqrt{\left(\mu-E_{j}\right)^{2}+\frac{1}{4\tau^{2}}}+\left(\mu-E_{j}\right)}}{\left(\mu-E_{j}\right)^{2}+\frac{1}{4\tau^{2}}}\left(1-\frac{2(\mu-E_{j})}{\sqrt{\left(\mu-E_{j}\right)^{2}+\frac{1}{4\tau^{2}}}}\right)}{\sum\limits_{j=1}^{J_{\max}}\sqrt{m_{j}}\frac{\sqrt{\sqrt{\left(\mu-E_{j}\right)^{2}+\frac{1}{4\tau^{2}}}+\left(\mu-E_{j}\right)}}{\sqrt{\left(\mu-E_{j}\right)^{2}+\frac{1}{4\tau^{2}}}}}.$ (21)
@@
Figure 4: The Seebeck coefficient according to Kelvin’s formula (a) and Mott’s formula (b) as a function of the chemical potential for the parabolic band with the quantum number j=1 for $\tau$=483.09 $eV^{-1}$ (blue line), $\tau$=38.65 $eV^{-1}$ (green line) and $\tau$=3.86 $eV^{-1}$ (red line) .
Effect of smearing of DOS in parabolic subbands.        The Mott’s formula. The smearing of the density of states jump in Eq. (12) will also affect the Mott’s formula for the Seebeck coefficient. Below we consider a relatively clean system and the range of not too low temperatures: $T\gg\hslash/\tau.$ Within the Lorentz approximation for delta function $\delta\left(E\right)=\frac{1}{\pi}\lim_{\tau\rightarrow\infty}\tau^{-1}/\left(E^{2}+\tau^{-2}\right)$ and using the relation $\delta\left(E\right)=\theta^{\prime}\left(E\right)$, one can present $\theta$-function in Eq. (12) in the form
$\theta\left(E\right)=\frac{1}{2}+\frac{1}{\pi}\arctan\left(E\tau\right).$ (22)

### v2

# Failure of Mott’s formula for the Thermopower in Carbon Nanotubes

@@ Abstract
Well-known Mott’s formula links the thermoelectric power characterised by Seebeck coefficient to conductivity. We calculate analytically the thermoelectric current and Seebeck coefficient in one-dimensional systems and show that, while the prediction of Mott’s formula is valid for Dirac fermions, it is misleading for the carriers having a parabolic dispersion. We apply the developed formalism to metallic single wall carbon nanotubes and obtain a non-trivial non-monotonic dependence of the Seebeck coefficient on the chemical potential. We emphasize that, in contrast to Mott’s formula, the classical Kelvin’s formula that links thermoelectric power to the temperature derivative of the chemical potential is perfectly valid in carbon nanotubes in the ballistic regime. Interestingly, however, the Kelvin’s formula fails in two- and three-dimensional systems in the ballistic regime.
@@ I Introduction
$S_{K}=\frac{1}{e}\left(\frac{\partial\mu}{\partial T}\right)_{N,V}$ (2)
(here $N$ is the number of particles, $V$ is the volume of the system).
The alternative, kinetic approach to the theoretical description of the Seebeck effect was developed by Sir Nevil Mott in the second half of the XXth century [9, 10]. The Mott’s formula relates the Seebeck coefficient to the conductivity of the system [11, 12]:
@@
where $\sigma$ is conductivity, $\mu$ is the chemical potential, $T$ is temperature. Here and further we assume $k_{B}=1.$ One can see that substituting the above expression for conductance to Mott’s formula one obtains $S_{M}=0$ for any value of chemical potential except the close vicinity of the bottoms of the quantisation subband. Below we demonstrate that this simple conclusion fails and Mott’s formula is unable to describe the thermoelectric phenomena in 1D systems in the ballistic regime. In contrast, the Kelvin’s formula remains valid in the ballistic regime in any one-dimensional (1D) system. Interestingly, however, the Kelvin’s formula turns out to be no more valid in two- and three-dimensional (2D and 3D) systems, in the ballistic regime. It is worth noting that the Seebeck coefficient can also be introduced through a more general expression for the conductivity within the relaxation time approximation [13]. However, we deliberately chose the original definition of Mott formula, which can be more easily extended to the case of the ballistic regime. We restrict our consideration to the ballistic, semi-classical, linear response regime. The thermocurrent and thermoelectricity in the ballistic regime have been carefully studied theoretically in 1D, 2D and 3D systems [14, 15, 16]. The difference between these works and our present work is that we focus on the broken circuit geometry where no electric current is flowing through the system. We present an analytical theory of the Seebeck effect in metallic CNTs in the regime of a ballistic transport. We note that the Seebeck coefficient is also a direct measure of the entropy per particle, which makes it one of the most important characteristics of the statistics of quasiparticles in crystals [17]. We find that, in CNTs, it is dramatically dependent on the type of electronic dispersion. In the case of a parabolic dispersion, the Seebeck coefficient is a non-monotonic function of the chemical potential. Its magnitude decreases with the increase of the chemical potential, and its sign changes in the vicinity of the resonances of the chemical potentials and bottom energies of the electron and hole quantization subbands (the effect of Seebeck coefficient sign change is well established experimental fact often observed in the vicinity of the Fermi surface topology changes [18, 19, 20], see also reviews [21, 22] and references therein.). In contrast, in the case of a linear dispersion that is observed in the vicinity of Dirac points in conducting CNTs the Seebeck coefficient is equal to zero. The sharp contrast between the behaviors of conductivity and Seebeck coefficient as functions of the chemical potential that may be efficiently controlled by an applied bias opens room for a variety of non-trivial effects governed by an interplay of currents induced by electric field and temperature gradient.
II Seebeck coefficient in the ballistic regime in a 1D system
II.1 1D Kelvin’s formula in the ballistic regime
The definition of the Seebeck coefficient in the case of a broken electric circuit (${\cal J}=0$, where ${\cal J}$ is an electric current) reads
@@ II Seebeck coefficient in the ballistic regime in a 1D system
Here and in what follows we assume the degeneracy factor of quantum states $g=1$. The results can be easily generalized for any other value of $g$. The electronic velocity is defined by the general relation:
@@
As expected, in every single band their product is constant $\nu(E)v(E)=\frac{1}{h}.$
Note that at negative energies, the Seebeck effect is dominated by holes rather than electrons. Holes are conveniently described by positive effective masses and, consequently, they have positive group velocities. The energy dependence of the hole contribution to the Seebeck effect is symmetric to one of the electron contribution taken with a negative sign as we discuss in Appendix B. We show, in particular, that at the Dirac point, electron and hole contributions to the Seebeck coefficient exactly compensate each other.
@@
Coming back to the 1D case, at low enough temperatures ($T\ll\mu$) one can rewrite the Kelvin’s formula in terms of the logarithmic derivative of DOS (See Appendix A).
$S_{K}=\frac{1}{e}\left(\frac{\partial\mu}{\partial T}\right)_{N,V}=\frac{\pi^{2}}{3e}T\frac{\partial}{\partial\mu}\ln\nu(\mu).$ (11)
II.2 1D Mott’s formula in the ballistic regime
It is instructive to compare the exact expression for the Seebeck coefficient obtained above with an approximation that follows from Mott’s formula. In order to derive it, we start from the generic expression for the electric current density in a one-dimensional channel in the ballistic regime (5).
@@
$\nu(E)=\frac{1}{{2\pi}}{\left|{\frac{{\partial E}}{{\partial k}}}\right|^{-1}}\left[{1+\theta\left({E-{E_{p}}}\right)}\right],$ (12)
@@
This expression can be now substituted to Mott’s formula (3). As a result we obtain:
@@
III The specific case of a metallic CNT
III.1 Energy spectra of single-walled carbon nanotubes
Figure 1: (a) A sketch of a zigzag CNT with chirality $(9,0)$. (b)The energy band structure of a metallic CNT according to the dispersion relation given by Eq. (15) in the limit of zero gap: $E_{0}=0$. Blue lines correspond to the energy levels with a quantum number $j=-1$ and $j=1$, green lines are for the energy levels with $j=0$.
Basic electronic properties of single-walled (dubbed originally as single-shell [23]) carbon nanotubes (CNTs) have been well-understood for over three decades [24] and depend entirely on their chirality $(m,n)$, where (as explained in detail in Ref. [25]) the numbers $(m,n)$ express the wrapping ‘chiral’ vector connecting two carbon sites along the nanotube circumference via the two unit vectors of the underlying hexagonal graphene lattice (see Fig. 1a). A CNT with $n-m\neq 3p$, where $p$ is an integer, is a quasi-one-dimensional semiconductor with the band gap $E_{g}$ inversely proportional to the nanotube radius $R=a\sqrt{n^{2}+m^{2}+nm}/(2\pi)\approx 0.39\sqrt{n^{2}+m^{2}+nm}$, where the graphene lattice constant $a=\sqrt{3}a_{C-C}$ with $a_{C-C}=1.42$ being a distance between neighbouring carbon atoms in graphene. Only armchair $(n,n)$ nanotubes are truly metallic (gapless) with the low-energy part of dispersion well-described by $E=\pm\hbar v_{F}|k-k_{0}|$, where $v_{F}\approx 9.8\times 10^{5}$ m/s is the Fermi velocity in graphene and $k_{0}=\pm 2\pi/(3a)$. The remaining CNTs with $n-m=3p$ and $n\neq m$, which are gapless within the frame of a simple zone-folding approximation [24, 25] of the $\pi$-electron graphene spectrum, are in fact narrow-gap semiconductors due to the curvature effects [26]. For illustration of our main results we use in this work parameters of a (9,0) zigzag nanotube sketched in Fig. 1(a) assuming it to be metallic with the curvature-induced band gap neglected. The energy spectrum of such a ”model” CNT is shown in Fig. 1(b).
@@ III.1 Energy spectra of single-walled carbon nanotubes
For the two branches closest to the Dirac point ($j=0$), for the metallic CNTs $E_{0}=0$ (see Fig. 1b). The bottoms of their higher subbands are given by $E_{j}=j\hbar v_{F}/R$. Notably, the subbands with $|j|\geq 1$ are reasonably-well described by the effective mass approximation with $m_{j}=E_{j}/v_{F}^{2}$.
III.2 A vanishing contribution to the Seebeck coefficient from linear dispersion subbands
For the linear dispersion subband DOS is constant and given by $\nu_{0}(\mu)=\frac{2}{\pi\hbar v_{F}}$. As a result, the temperature derivative of the concentration of charge carriers turns zero. Hence the Dirac mode does not contribute to the Seebeck coefficient in agreement with both the Kelvin’s and Mott’s formulae. This statement remains valid also in the vicinity of the Dirac point due to the compensation of electron and hole contributions to the Seebeck effect (see Appendix B).
III.3 Conflicting predictions of the Kelvin’s and Mott’s approaches for the contributions of parabolic subbands to the Seebeck coefficient
Figure 2: (a) The conductance (blue line) for $T=2.4$ K given by Eq. (13) and the DOS of a metallic CNT without the smearing effect (inset) given by Eq. (7). Here $\nu_{0}$ is a density of states in the central linear subband and $\sigma_{0}$ is its contribution to the conductance. (b) DOS of a metallic CNT without the smearing effect (black line) and with smearing one for $\tau/\hbar$=483.09 $eV^{-1}$ (blue line), $\tau/\hbar$=38.65 $eV^{-1}$ (green line) and $\tau/\hbar$=3.86 $eV^{-1}$ (red line). The vertical black arrow at the top of (b) specify the energy corresponding to the bottom of the first subband with $j=1$. Other parameters of a metallic CNT are ${v_{F}}=9.8\cdot{10^{5}}\rm{m}/\rm{s}$, ${a_{c}}=0.142\,{\rm{nm}}$ and chirality $(9,0)$, which gives $R=0.3522\,{\rm{nm}}$.
Figure 2(a) shows the conductance calculated using Eq. (13) and the density of states (DOS) calculated with Eq. (7) for a metallic CNT characterised with the dispersion of electronic bands shown in Fig. 1(b). The comparison of these curves helps understanding of the qualitative difference between the Seebeck coefficients obtained with the Kelvin’s and the Mott’s formula governed by the derivatives of the DOS and conductivity logarithms over the chemical potential, respectively. Indeed, while DOS in a one-dimensional system is a non-monotonic function of energy, the conductivity is a monotonic staircase-like function. This is why, according to the Kelvin’s formula the Seebeck coefficient changes its sign at each size quantization subband bottom, while according to the Mott’s formula, it remains always positive.
It is also important to note that in all numerical calculations presented in this work we refer to a metallic CNT whose energy spectrum is presented in Fig. 1(b). It contains of the first, linear, band shown by a green line in Fig. 1(b) and the second, parabolic, band shown by the blue line in Fig. 1(b). As we demonstrate analytically, the contribution to the Seebeck coefficient from the linear band is negligibly small at low temperatures. Only the second, parabolic band (and upper bands that are not considered here) gives a sizeable contribution to the Seebeck coefficient. In this context, while one always needs to sum over subbands as it is done in Eq. (17), for the specific considered case of a metallic CNT, the single band contribution in Eq. (14) is dominant.
@@ III.2 A vanishing contribution to the Seebeck coefficient from linear dispersion subbands
$\nu(\mu)=\sum\limits_{j=1}^{J_{\max}}\frac{\sqrt{2m_{j}}}{\pi\hbar}\frac{1}{\sqrt{\mu-E_{j}}},$ (16)
@@
Figure 3: The Seebeck coefficient for a metallic CNT as a function of chemical potential $\mu$ calculated for three temperatures T = 2.4 K (a), T = 77 K (b) and T = 300 K (c) according to Kelvin’s formula Eq. (17) (red lines) and Mott’s formula Eq. (14) (blue lines). The vertical black arrows at the top of the figures specify the energy corresponding to the bottom of the first subband with $j=1$.
@@
The dependence of the Seebeck coefficient on the chemical potential calculated at different temperatures is shown in Fig. 3. One can notice the strong difference in the results obtained following the Kelvin’s approach (red lines) and Mott’s approach (blue lines). The Kelvin’s approach predicts the hyperbolic decay of the Seebeck coefficient away from the resonance between the chemical potential and the bottom of the subband. At the resonance point, the singularity and the change of sign of the Seebeck coefficient are observed. In contrast, Mott’s formula predicts a finite value of the Seebeck coefficient at the bottom of subband, no sign change and a fast exponential decay away from the resonance. We note at this point that long ago the change of the sign of the Seebeck coefficient was identified as a signature of the topological phase transition which indeed takes place ones the chemical potential crosses the bottom of the next quantization subband [18, 27, 19, 21, 28, 29, 30]. The divergence of the Seebeck coefficient in the absence of scattering is caused by the singularity in DOS. It disappears once the smearing of the density of states is taken into account as we show in the next section.
IV Effect of smearing of DOS in parabolic Subbands
IV.1 A parabolic subband: the Kelvin’s formula
In any realistic 1D system including CNTs, singularities of DOS are smeared due to a variety of factors: from finite size effects to fluctuations and non-linearities. Therefore, it is important to consider an effect of smearing of DOS on the Seebeck coefficient in the ballistic regime. The Green’s function of an electron in a single-wall CNT is
$G_{j}^{R}(E,p)=\frac{1}{E-p_{z}^{2}/2m_{j}-E_{j}+i\hbar/2\tau},$ (18)
where $\tau$ is the positive parameter responsible for the smearing effect and $p_{z}$ is the continuous momentum along the $z$-direction. This smearing is the result of electron-phonon interaction.
@@ IV.1 A parabolic subband: the Kelvin’s formula
$\nu^{(sm)}(E)=\sum\limits_{j=1}^{J_{\max}}\frac{\sqrt{m_{j}}}{\pi\hbar}\frac{\sqrt{\sqrt{\left(E-E_{j}\right)^{2}+\frac{\hbar^{2}}{4\tau^{2}}}+\left(E-E_{j}\right)}}{\sqrt{\left(E-E_{j}\right)^{2}+\frac{\hbar^{2}}{4\tau^{2}}}},$ (20)
@@
Now one can derive the contribution to the Seebeck coefficient of a CNT coming from parabolic energy subbands:
$S^{(sm)}_{K}(\mu)=\frac{\pi^{2}T}{3eJ_{\max}}\frac{\partial\ln\nu^{(sm)}(\mu)}{\partial\mu}=\frac{\pi^{2}T}{6eJ_{\max}}\frac{\sum\limits_{j=1}^{J_{\max}}\sqrt{m_{j}}\frac{\sqrt{\sqrt{\left(\mu-E_{j}\right)^{2}+\frac{\hbar^{2}}{4\tau^{2}}}+\left(\mu-E_{j}\right)}}{\left(\mu-E_{j}\right)^{2}+\frac{\hbar^{2}}{4\tau^{2}}}\left(1-\frac{2(\mu-E_{j})}{\sqrt{\left(\mu-E_{j}\right)^{2}+\frac{\hbar^{2}}{4\tau^{2}}}}\right)}{\sum\limits_{j=1}^{J_{\max}}\sqrt{m_{j}}\frac{\sqrt{\sqrt{\left(\mu-E_{j}\right)^{2}+\frac{\hbar^{2}}{4\tau^{2}}}+\left(\mu-E_{j}\right)}}{\sqrt{\left(\mu-E_{j}\right)^{2}+\frac{\hbar^{2}}{4\tau^{2}}}}}.$ (21)
@@
Figure 4: The Seebeck coefficient according to Kelvin’s formula (a) and Mott’s formula (b) as a function of the chemical potential for the parabolic band with the quantum number j=1 for $\tau/\hbar$=483.09 $eV^{-1}$ (blue line), $\tau/\hbar$=38.65 $eV^{-1}$ (green line) and $\tau/\hbar$=3.86 $eV^{-1}$ (red line) .
IV.2 A parabolic subband: Mott’s formula
The smearing of the density of states jump in Eq. (12) will also affect Mott’s formula for the Seebeck coefficient. Below we consider a relatively clean system and the range of not too low temperatures: $T\gg\hslash/\tau.$ Within the Lorentz approximation for delta function $\delta\left(E\right)=\frac{1}{\pi}\lim_{\tau\rightarrow\infty}\hbar\tau^{-1}/\left(E^{2}+\hbar^{2}\tau^{-2}\right)$ and using the relation $\delta\left(E\right)=\theta^{\prime}\left(E\right)$, one can present $\theta$-function in Eq. (12) in the form
$\theta\left(E\right)=\frac{1}{2}+\frac{1}{\pi}\arctan\left(E\hbar^{-1}\tau\right).$ (22)